The Walton Company is expected to pay a $4 dividend. Over the next 12 months, the dividend is expected to grow at 5 percent, which is the constant growth rate for the firm. Using an 12% required rate of return, compute the price of a common share.
1
Expert's answer
2020-04-23T18:25:46-0400

The price of a common share is:

"P = D0*\frac{1 + g}{r - g} = 4*\frac{1 + 0.05}{0.12 - 0.05} =" $60.